OpenStudy (anonymous):

Your friend hands you a graph of the performance statistics of the newest car to roll off the assembly line. He says, “I know this graph is f(x) = –4(x – 5)3 + 7 but I can’t remember how it is related to the graph of x3.” Explain to your friend how the graph f(x) is a translation of the graph x3.

OpenStudy (anonymous):

this one is hard i had it yesterday. Are you on flvs

OpenStudy (anonymous):

Yeah @andy8150

OpenStudy (anonymous):

what teacher do you have

OpenStudy (anonymous):

niedzwiecki

OpenStudy (anonymous):

ohh

OpenStudy (anonymous):

Can you help or nah?

OpenStudy (anonymous):

here is a graph

OpenStudy (anonymous):

thats all i know

OpenStudy (anonymous):

Thanks anyways

OpenStudy (anonymous):

@hartnn could you help?

OpenStudy (anonymous):

@Nurali

OpenStudy (anonymous):

@zepdrix

zepdrix (zepdrix):

\[\Large\rm y=(x)^3\]Let's try to turn this into the performance function. Putting a negative in front flips it over the x-axis.\[\Large\rm y=-(x)^3\]To make shifts to the `right` or `up` we subtract that value from the corresponding variable. So we want to shift this function to the `right` 5 units, so we'll subtract 5 from x.\[\Large\rm y=-(x-5)^3\]We also want to shift this function `up` by 7, so we'll subtract 7 from y.\[\Large\rm y-7=-(x-5)^3\]Then add 7 to each side,\[\Large\rm y=-(x-5)^3+7\]Then we'll `squeeze` the function (make it grow four times faster) by multiplying it by 4.\[\Large\rm y=-4(x-5)^3+7\]

zepdrix (zepdrix):

Hopefully that helps to explain some of those changes :)